OpenStudy (anonymous):

HELP! Award At End :) I am stuck and no idea.. Given r(t), find the following...r(t) = < t^2, sin(t) - tcos(t), cos(t) + tsin(t)> Find the length of the curve from 0 to pi.

OpenStudy (loser66):

we have the formula to solve the problem r(t) = < f(t), g(t), h(t)> \(a\leq t\leq b\) then arclength \(L = \int_a^b \sqrt{f'(t)^2+g'(t)^2+h'(t)^2}dt\) so, just apply

OpenStudy (anonymous):

So t^2 is 2t. sin(t) - tcos(t) is 2cos(t) - tsin(t). cos(t) + tsin(t) is -tcost Then I just put that into the arcLength Formula. Take the square root of it, and then set t = 0 and pi?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

f'(t) = 2t but \(f'(t)^2= 4t^2\)

OpenStudy (loser66):

and you take derivative of g(t) wrong. Redo, please